Natalie has been approached by Ken Thornton, a shareholder of The Beanery Coffee Inc. Ken wants to retire and would like to sell his 1,000 shares in The Beanery Coffee, which represents 30% of all shares issued. The Beanery is currently operated by Ken’s twin daughters, who each own 35% of the common shares. The Beanery not only operates a coffee shop but also roasts and sells beans to retailers, under the name “Rocky Mountain Beanery.”

The business has been operating for approximately 5 years. In the last 2 years Ken has lost interest and left the day-to-day operations to his daughters. Both daughters at times find the work at the coffee shop overwhelming. They would like to have a third shareholder involved to take over some of the responsibilities of running a small business. Both feel that Natalie and Curtis are entrepreneurial in spirit and that their expertise would be a welcome addition to the business operation. The twins have also said that they plan to operate this business for another 10 years and then retire.

Ken has met with Curtis and Natalie to discuss the business operation. All have concluded that there would be many advantages for Cookie & Coffee Creations Inc. to acquire an interest in The Beanery Coffee. One of the major advantages would be volume discounts for purchases of the coffee bean inventory.

Despite the apparent advantages, however, Natalie and Curtis are still not convinced that they should participate in this business venture.

Assume that Ken wants to sell his 1,000 shares of The Beanery Coffee for $15,000.

Prepare the journal entry required if Cookie & Coffee Creations Inc. buys Ken’s shares. (If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Account Titles and Explanation

Debit

Credit

Assume that Cookie & Coffee Creations Inc. buys the shares and in the following year The Beanery Coffee earns $50,000 net income and pays $25,000 in dividends.

Prepare the journal entries required under both the cost method and the equity method of accounting for this investment. (If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Account Titles and Explanation

Debit

Credit

Cost Method:

Equity Method:

(To record equity in Beanery’s net income)

(To record dividends received)

Identify where this investment would be classified on the balance sheet of Cookie & Coffee Creations Inc.
Investment would be classified as a

Current AssetsCurrent LiabilitiesExpensesIntangible AssetsLong-term InvestmentsLong-term LiabilitiesNet Income / (Loss)Property, Plant and EquipmentRevenuesStockholders' Equity

in balance sheet.
What amount would appear on the balance sheet under each of the methods of accounting for the investment?
Cost method $
Equity method $
Answer #1

1. Journal Entries:

By using Cost method:

Particulars Debit Credit
Investment in stocks of beanery coffee Inc. 15,000
Cash        15,000
(Being shares purchased)
Cash 7,500
Dividend Income 7,500
(Being dividend received)

2. Journal entries:

By using equity method:

Investment in stocks of beanery coffee Inc. = 50,000 * 30% = $15,000

Investment in stocks of beanery coffee Inc. = 25,000 * 30% = $7,500

Particulars Debit Credit
Investment in stocks of beanery coffee Inc. 15,000
Cash        15,000
(Being shares purchased)
Investment in stocks of beanery coffee Inc. 15,000
Income from Investment in beanery company 15,000
(Being income earned)
Cash 7,500
Investment in stocks of beanery coffee Inc. 7,500
(Being dividend received)

3. Investment would be classified as a Long term investment in balance sheet of cookie & coffee creations Inc.

4. Investment amount appears on the balance sheet under:

Cost method = Cost of acquisition

= $15,000

Therefore, Investment amount appear on the balance sheet under cost method = $15,000.

Investment amount appears on the balance sheet under:

Equity method = Cost of acquisition + income received - Dividend received

= 15,000 + 15,000 - 7,500

= $22,500

Therefore, Investment amount appear on the balance sheet under Equity method = $22,500